# Spectral gap for surfaces of infinite volume with negative curvature

### v1

@@ Abstract.
We prove that the imaginary parts of scattering resonances for negatively curved asymptotically hyperbolic surfaces are uniformly bounded away from zero and provide a resolvent bound in the resulting resonance-free strip. This provides an essential spectral gap without the pressure condition. This is done by adapting the methods of [NSZ11], [Vas13a] and [Vac22] and answers a question posed in [DyZa16].
@@ 1. Introduction
In a seminal paper Bourgain–Dyatlov [BoDy18] showed that a convex cocompact hyperbolic surface has an essential spectral gap between the unitarity axis and the set of scattering resonances. This means the Selberg zeta function has only finitely many zeros in the region $\mathop{\rm Re}\nolimits s>1/2-\beta$ for some $\beta>0$. This holds without any assumptions on the Hausdorff dimension of the trapped set, in particular without a “pressure condition” which in this case goes back to the works of Patterson and Sullivan [Pa76, Su79]. The purpose of this note is to generalize this result to negatively curved surfaces which are asymptotically hyperbolic in a sense described below. This is done by combining the quantum monodromy method of Nonnenmacher–Sjöstrand–Zworski [NSZ11] and Vasy’s method for meromorphic continuation [Vas13a, Zw16, DyZw19] with the recent work of Vacossin [Vac22]. It answers a question posed by Dyatlov–Zahl [DyZa16] in the first paper on the fractal uncertainty principle.
@@
The proof of the main Theorem follows from [Vac22] by reducing the problem to quantum monodromy maps using [NSZ11]. Although [Vac22] also uses [NSZ11], our approach is different by replacing the application of microlocal weight functions with propagation estimates. This approach simplifies some aspects of [NSZ11] and allows a seamless application of Vasy’s method [Vas13a]. The geometric component comes from the now classical work of Eberlein [Eb72] which shows the trapped set is topologically one dimensional in our setting (see §2.3).
The spectral gap for open hyperbolic quantum systems has been studied since Ikawa [Ik88] in mathematics and Gaspard–Rice [GaRi89] in physics – if the topological pressure (an object from thermodynamical formalism) satisfies $P(1/2)<0$, the statement of the theorem above holds with $\beta<-P(1/2)$. For an experimental manifestation of this gap, see [B13]. A general spectral gap under the pressure condition was proved by Nonnenmacher–Zworski [NoZw09a, NoZw09b]. The first advances in the direction of improving the pressure gaps were made by Naud [Na05] (in the setting of constant curvature surfaces and complex dynamics) and Petkov–Stoyanov [PeSt10] (in the setting of obstacle scattering). These results were based on Dolgopyat’s method [Do98]. However, it was conjectured by Zworski [Zw17, Conjecture 3] that the pressure condition is not necessary. Dyatlov–Zahl [DyZa16] made the first breakthrough showing a spectral gap without the pressure condition by introducing the fractal uncertainty principle. Bourgain–Dyatlov [BoDy18] proved the fractal uncertainty principle for any porous set in dimension one and showed that any (noncompact) convex cocompact hyperbolic surface has an essential spectral gap. For recent advances on the fractal uncertainty principle, see [BLT23, Co23]. The spectral gap was generalized to open quantum maps in dimension $2$ by Vacossin [Vac22, Vac23] by combining the method of [BoDy18] and [DJN22]. That provided an essential spectral gap for classes of obstacles and semiclassical scattering problems.
@@ Acknowledgement
We thank Maciej Zworski for many helpful discussions, and for his encouragement to write this note. We thank Stéphane Nonnenmacher for helpful discussions on Vacossin’s paper [Vac22]. The author was partially supported by the NSF grant DMS-1952939 and by the Simons Targeted Grant Award No. 896630.
@@ 2.1. Review of concepts
${\mathcal{F}}_{h}(\chi u)(\xi)={\mathcal{O}}(h^{\infty}\langle\xi\rangle^{-\infty}),\quad\chi\in C_{c}^{\infty}(U),\xi\in V\cap{\mathbb{R}}_{\xi}^{d}.$
The semiclassical wavefront set of an operator $A:C_{c}^{\infty}(Y)\to{\mathcal{D}}^{\prime}(X)$ with $h$-tempered Schwartz kernel ${\mathcal{K}}_{A}$ is defined as (see [DyZw19, Definition E.36])
@@ 2.2. Vasy’s method revisited
    is a holomorphic Fredholm family of index 00.
@@
$\Gamma_{\pm}:=\{(x,\xi)\in T^{*}X\setminus 0:\exp(tH_{p})(x,\xi)\text{ remainsbounded as }t\to\mp\infty\}\cap p^{-1}(0).$
@@
We remark that similar modifications are also used in [JiTa23, Proposition 2.7].
@@ 3. Quantum monodromy without weights
In order to apply [Vac22] to obtain the spectral gap, we need to construct quantum monodromy maps as in [NSZ11] but in the asymptotically hyperbolic setting. In §3.1 we apply Vasy’s method to construct such quantum monodromy maps without using weights for $d+1$-dimensional asymptotically hyperbolic manifolds whose trapped set has topological dimension $1$. In §3.2 we use [Eb72] to verify this assumption for all asymptotically hyperbolic surfaces.
@@ 3.1.1. Dynamical preliminaries
$t_{0}:=\min\{t>0:\text{ there exist }j\neq k\text{ and }(x,\xi)\in\Sigma_{k}\text{ such that }\varphi^{t}(x,\xi)\in\Sigma_{j}\}>0.$ (3.1)
@@ 3.1.2. Microlocal Grushin problem
$\operatorname{WF}_{h}^{\prime}(\widetilde{E})$$\subset\{(\varphi^{t}(x,\xi);x,\xi):t>-\epsilon\},$ (3.3)
@@ 3.1.3. Global Grushin problem
Since ${\mathcal{P}}(z)$ is a Fredholm operator of index 00, it suffices to construct a right inverse of ${\mathcal{P}}(z)$. In other words, given $(v,v_{+})\in\bar{H}_{h}^{s-1}({X_{1}})\times{\mathcal{H}}_{h}$ with $\|v\|_{\bar{H}_{h}^{s-1}}^{2}+\|v_{+}\|_{L^{2}}^{2}\leq 1$, we want to find $(u,u_{-})\in D_{h}^{s}({X_{1}})\times{\mathcal{H}}_{h}$ such that
@@
we aime to solve for ${\mathcal{P}}(z)(u,u_{-})^{T}=(v_{0},v_{+}-v_{0+})^{T}$. At this point we can use the microlocal Grushin problem to find $(\widetilde{u},\widetilde{u}_{-})\in H(V_{2})\times H(\widetilde{D}\cup\widetilde{A})$ such that
@@
Suppose $\text{Suppose }U\text{ is a neighbourhood of the trapped set }K_{0}.$ (3.8)
Then there exists a smaller neighbourhood $\text{Then there exists a smaller }\text{neighbourhood }U_{0}\subset U\text{ such that}$
$\text{if }(x,\xi)\in U_{0}\text{ and }\varphi^{t}(x,\xi)\notin U\text{ for some }t>0,\text{ then }(x,\xi)\notin\Gamma_{-}.$
@@
Eventually $\operatorname{WF}_{h}({\mathcal{R}}^{N}(v,v_{+})^{T})\cap\operatorname{WF}_{h}(\widetilde{R}_{-})=\varnothing$ and thus there exists $N_{1}\in\mathbb{N}$ with ${\mathcal{R}}^{N_{1}}=\mathcal{O}(h^{\infty})$. Let
@@ 4. Proof of the spectral gap and the resolvent estimate
We have reduced the problem to quantum monodromy maps in §3.1. In the case of surfaces, [Vac22, Proposition 4.1] shows

### v2

@@ Abstract.
We prove that the imaginary parts of scattering resonances for negatively curved asymptotically hyperbolic surfaces are uniformly bounded away from zero and provide a resolvent bound in the resulting resonance-free strip. This provides an essential spectral gap without the pressure condition. This is done by adapting the methods of [NSZ11], [Vas13a], and [Vac24] and answers a question posed in [DyZa16].
@@ 1. Introduction
In a seminal paper, Bourgain–Dyatlov [BoDy18] showed that a convex cocompact hyperbolic surface has an essential spectral gap between the unitarity axis and the set of scattering resonances. This means the Selberg zeta function has only finitely many zeros in the region $\mathop{\rm Re}\nolimits s>1/2-\beta$ for some $\beta>0$. This holds without any assumptions on the Hausdorff dimension of the trapped set, in particular without a “pressure condition” which in this case goes back to the works of Patterson and Sullivan [Pa76, Su79]. The purpose of this note is to generalize this result to negatively curved surfaces which are asymptotically hyperbolic in a sense described below. This is done by combining the quantum monodromy method of Nonnenmacher–Sjöstrand–Zworski [NSZ11] and Vasy’s method for meromorphic continuation [Vas13a, Zw16, DyZw19] with the recent work of Vacossin [Vac24]. It answers a question posed by Dyatlov–Zahl [DyZa16] in the first paper on the fractal uncertainty principle.
@@
The proof of the main Theorem follows from [Vac24] by reducing the problem to quantum monodromy maps using [NSZ11]. Although [Vac24] also uses [NSZ11], our approach is different by replacing the application of microlocal weight functions with propagation estimates. This approach simplifies some aspects of [NSZ11] and allows a seamless application of Vasy’s method [Vas13a]. The geometric component comes from the now classical work of Eberlein [Eb72] which shows the trapped set is topologically one dimensional in our setting (see §2.3).
The spectral gap for open hyperbolic quantum systems has been studied since Ikawa [Ik88] in mathematics and Gaspard–Rice [GaRi89] in physics – if the topological pressure (an object from thermodynamical formalism) satisfies $P(1/2)<0$, the statement of the theorem above holds with $\beta<-P(1/2)$. For an experimental manifestation of this gap, see [B13]. A general spectral gap under the pressure condition was proved by Nonnenmacher–Zworski [NoZw09a, NoZw09b]. The first advances in the direction of improving the pressure gaps were made by Naud [Na05] (in the setting of constant curvature surfaces and complex dynamics) and Petkov–Stoyanov [PeSt10] (in the setting of obstacle scattering). These results were based on Dolgopyat’s method [Do98]. However, it was conjectured by Zworski [Zw17, Conjecture 3] that the pressure condition is not necessary. Dyatlov–Zahl [DyZa16] made the first breakthrough showing a spectral gap without the pressure condition by introducing the fractal uncertainty principle. Bourgain–Dyatlov [BoDy18] proved the fractal uncertainty principle for any porous set in dimension one and showed that any (noncompact) convex cocompact hyperbolic surface has an essential spectral gap. For recent advances on the fractal uncertainty principle, see [BLT25, Co25a, Co25b]. The spectral gap was generalized to open quantum maps in dimension $2$ by Vacossin [Vac24, Vac23] by combining the method of [BoDy18] and [DJN22]. That provided an essential spectral gap for classes of obstacles and semiclassical scattering problems.
The bound (1.3) was recently used in Huang–Sogge–Tao–Zhang [HSTZ25] to show lossless Strichartz estimates and spectral projection estimates in the same setting.
@@ Acknowledgement
We thank Maciej Zworski for many helpful discussions, and for his encouragement to write this note. We thank Stéphane Nonnenmacher for helpful discussions on Vacossin’s paper [Vac24]. We thank Semyon Dyatlov for helpful discussions on the appendix, especially for explaining the compactness argument to me. We also thank Jialun Li and Wenyu Pan for the example of rigid acylindrical hyperbolic manifolds. The author was partially supported by the NSF grant DMS-1952939 and by the Simons Targeted Grant Award No. 896630.
@@ 2.1. Review of concepts
${\mathcal{F}}_{h}(\chi u)(\xi)={\mathcal{O}}(h^{\infty}\langle\xi\rangle^{-\infty}),\quad\chi\in C_{c}^{\infty}(U),\xi\in V\cap{\mathbb{R}}_{\xi}^{d},$
where
${\mathcal{F}}_{h}u(\xi):=\frac{1}{(2\pi h)^{d/2}}\int_{{\mathbb{R}}^{d}}u(x)e^{-ix\cdot\xi/h}dx$ (2.1)
is the semiclassical Fourier transform. The semiclassical wavefront set of an operator $A:C_{c}^{\infty}(Y)\to{\mathcal{D}}^{\prime}(X)$ with $h$-tempered Schwartz kernel ${\mathcal{K}}_{A}$ is defined as (see [DyZw19, Definition E.36])
@@ 2.2. Vasy’s method revisited
    is a holomorphic Fredholm family of index 0.
@@
$\Gamma_{\pm}:=\{(x,\xi)\in T^{*}X\setminus 0:\exp(tH_{p})(x,\xi)\text{ remains bounded as }t\to\mp\infty\}\cap p^{-1}(0).$ (2.7)
@@
We remark that similar modifications are also used in [JiTa25, Proposition 2.7].
@@ 3. Quantum monodromy without weights
In order to apply [Vac24] to obtain the spectral gap, we need to construct quantum monodromy maps as in [NSZ11] but in the asymptotically hyperbolic setting. In §3.1 we apply Vasy’s method to construct such quantum monodromy maps without using weights for $d+1$-dimensional asymptotically hyperbolic manifolds whose trapped set has topological dimension $1$. In §3.2 we use [Eb72] to verify this assumption for all asymptotically hyperbolic surfaces.
@@ 3.1.1. Dynamical preliminaries
[…]
@@ 3.1.2. Microlocal Grushin problem
[…]
@@ 3.1.3. Global Grushin problem
Since ${\mathcal{P}}(z)$ is a Fredholm operator of index 0, it suffices to construct a right inverse of ${\mathcal{P}}(z)$. In other words, given $(v,v_{+})\in\bar{H}_{h}^{s-1}({X_{1}})\times{\mathcal{H}}_{h}$ with $\|v\|_{\bar{H}_{h}^{s-1}}^{2}+\|v_{+}\|_{L^{2}}^{2}\leq 1$, we want to find $(u,u_{-})\in D_{h}^{s}({X_{1}})\times{\mathcal{H}}_{h}$ such that
@@
[…]
@@
[…]
@@
[…]
@@ 4. Proof of the spectral gap and the resolvent estimate
We have reduced the problem to quantum monodromy maps in §3.1. In the case of surfaces, [Vac24, Proposition 4.1] shows
@@
Appendix A Spectral gap for convex cocompact hyperbolic manifolds
In this appendix, we discuss a special case of the spectral gap problem in higher dimensions – convex cocompact hyperbolic manifolds, following Cohen’s recent breakthrough [Co25b]. We prove the following
Theorem A.1.
[…]
$R(\lambda)=(-\Delta-d^{2}/4-\lambda^{2})^{-1}:L^{2}_{\rm comp}(\Gamma\backslash\mathbb{H}^{d+1})\to L^{2}_{\rm loc}(\Gamma\backslash\mathbb{H}^{d+1})$
continues holomorphically from $\mathop{\rm Im}\nolimits\lambda>d/2$ to $\{|\lambda|>C_{0},\,\mathop{\rm Im}\nolimits\lambda>-\beta\}$, and we have the following resolvent estimate for any $\epsilon>0$:
$\|\chi(-\Delta-d^{2}/4-\lambda^{2})^{-1}\chi\|_{L^{2}\to L^{2}}\leq C_{\chi,\epsilon}|\lambda|^{-1-2\min(0,\mathop{\rm Im}\nolimits\lambda)+\epsilon},\quad|\lambda|>C_{0},\,\,\mathop{\rm Im}\nolimits\lambda>-\beta.$ (A.1)
Here $\chi\in C_{c}^{\infty}(\Gamma\backslash\mathbb{H}^{d+1})$ is a compactly supported cutoff.
Note that there are non-trivial examples of limit sets that contain a round circle, see e.g. [MMO16]. It is an open problem whether such manifolds have an essential spectral gap.
We expect the method of this paper will be helpful to further studies of the variable curvature case for asymptotically hyperbolic manifolds in higher dimensions.
Jin–Zhang–Zhang [JZZ26] recently gave a quantitative bound in Theorem A.4 below and an explicit essential spectral gap for quasi-Fuchsian groups.
A.1. Convex cocompact hyperbolic manifolds
[…]
The hyperbolic space $\mathbb{H}^{d+1}$ has (orientation preserving) isometry group $G=SO(d+1,1)_{0}$. This can be seen by identifying $\mathbb{H}^{d+1}$ with the sheet $x_{0}>0$ of
$\langle x,x\rangle_{\eta}=x_{1}^{2}+\cdots+x_{d+1}^{2}-x_{0}^{2}=-1$ (A.2)
in the Minkowski space ${\mathbb{R}}^{d+1,1}$ with metric $\eta=\sum_{1\leq j\leq d+1}dx_{j}^{2}-dx_{0}^{2}$. Let $\Gamma\subset G$ be a discrete torsion-free subgroup acting on $\mathbb{H}^{d+1}$ (strictly speaking, one needs to consider the non-orientable case; but we can always pass to an orientable double cover). We compactify $\mathbb{H}^{d+1}$ using the ball model (with the metric blowing up near the boundary). The limit set $\Lambda_{\Gamma}\subset\partial\mathbb{H}^{d+1}={\mathbb{S}}^{d}$ is defined as the set of limit points of a $\Gamma$-orbit of a point $\mathfrak{o}\in\mathbb{H}^{d+1}$. The convex core is defined as
[…]
$\widetilde{K}_{0}=\{(x,\xi)\in S^{*}\mathbb{H}^{d+1}:e^{tH_{p}}(x,\xi)\to\Lambda_{\Gamma},\,\,t\to\pm\infty\}\simeq(\Lambda_{\Gamma}\times\Lambda_{\Gamma}\setminus\Delta_{\Lambda_{\Gamma}})\times{\mathbb{R}}$
where $\Delta_{\Lambda_{\Gamma}}$ is the diagonal.
[…]
  1. (1)
    The geodesic vector field
    $Y=\begin{pmatrix}0&\cdots&0&0\\ \vdots&\ddots&\vdots&\vdots\\ 0&\cdots&0&1\\ 0&\cdots&1&0\end{pmatrix}$.
  2. (2)
    The horocyclic vector field
    $U_{j}^{\pm}=\begin{pmatrix}0&\cdots&0&\cdots&0&0\\ \vdots&\ddots&\vdots&\ddots&\vdots&\vdots\\ 0&\cdots&0&\cdots&\pm 1&-1\\ \vdots&\ddots&\vdots&\ddots&\vdots&\vdots\\ 0&\cdots&\mp 1&\cdots&0&0\\ 0&\cdots&-1&\cdots&0&0\end{pmatrix},\quad j=1,2,\cdots,d,$
    where the nonzero entries are in the (d+1j)(d+1-j)-th row and (d+1j)(d+1-j)-th column.
We have
$Y,U_{j}^{\pm}]=\pm U_{j}^{\pm},\quad[U_{j}^{+},U_{j}^{-}]=2Y.$ (A.3)
Note that $Y$ commutes with rotational vector fields in $\mathfrak{so}(d)$ and consequently descends to $S^{*}X$. However, $U_{j}^{\pm}$ does not commute with the rotational vector fields in $\mathfrak{so}(d)$ and does not descend to $S^{*}X$. Since
$U_{i}^{\pm},U_{j}^{\pm}]=0,\quad[U_{j}^{\pm},\mathfrak{so}(d)]\in\mathrm{span}(U^{\pm}),\quad[U_{i}^{\pm},U_{j}^{\mp}]\in\mathfrak{so}(d),\quad i\neq j,$
$\mathrm{span}(U^{\pm})$ is invariant under $SO(d)$ and thus descends to $S^{*}X$.
By (A.3), the stable bundle $E_{s}$ is spanned by projections of $U_{j}^{+}$ and the unstable bundle $E_{u}$ is spanned by projections of $U_{j}^{-}$. Let $\pi_{S}:F^{*}X\to S^{*}X$ be the projection, we can also write
$\pi_{S}^{*}(E_{s/u})=\mathrm{span}(U^{\pm})\oplus\mathfrak{so}(d).$
Now we write down the flows in coordinates. In the hyperboloid model (A.2), the sphere bundle $S^{*}\mathbb{H}^{d+1}$ can be identified with
$\{(x,\xi)\in{\mathbb{R}}^{d+1,1}:\langle x,x\rangle_{\eta}=-1,\,\,\langle x,\xi\rangle_{\eta}=0,\,\,\langle\xi,\xi\rangle_{\eta}=1,\,\,x_{0}>0\}$
and the frame bundle $F^{*}\mathbb{H}^{d+1}$ can be identifies as
$\{(x,\xi^{1},\cdots,\xi^{d+1}):x\in\mathbb{H}^{d+1},\,\,\xi^{1},\cdots,\xi^{d+1}\in T^{*}\mathbb{H}^{d+1}\text{ is a positively oriented orthonormal basis}\}.$
The map $\pi_{S}$ can be identified with $(x,\xi^{1},\cdots,\xi^{d+1})\to(x,\xi^{1})$. The geodesic flow $e^{tY}$ acts on $S^{*}\mathbb{H}^{d+1}$ by
$e^{tY}(x,\xi)=(x\cosh t+\xi\sinh t,x\sinh t+\xi\cosh t)$
and acts on $F^{*}\mathbb{H}^{d+1}$ by
$e^{tY}(x,\xi^{1},\cdots,\xi^{d+1})=(x\cosh t+\xi^{1}\sinh t,x\sinh t+\xi^{1}\cosh t,\xi^{2},\cdots,\xi^{d+1}).$
The horocyclic vector field $U_{j}^{\pm}$ acts on $F^{*}\mathbb{H}^{d+1}$ by (here the matrices act on the right)
$\begin{split}&e^{sU_{j}^{\pm}}(x,\xi^{1},\cdots,\xi^{d+1})\\ =&\left(x-s\xi^{j+1}+\frac{s^{2}}{2}(x\pm\xi^{1}),\xi^{1}\pm s\xi^{j+1}-\frac{s^{2}}{2}(\xi^{1}\pm x),\xi^{2},\cdots,\xi^{j+1}\mp s\xi^{1}-sx,\cdots,\xi^{d+1}\right).\end{split}$
Under the Hopf parametrization (we denote $x=(x^{\prime},x_{0})$ and $\xi=(\xi^{\prime},\xi_{0})$)
$\begin{split}\Phi:S^{*}\mathbb{H}^{d+1}\simeq({\mathbb{S}}_{+}^{d}\times{\mathbb{S}}_{-}^{d}\setminus\Delta_{{\mathbb{S}}^{d}})\times{\mathbb{R}}\\ (x,\xi)\to\left(\frac{x^{\prime}+\xi^{\prime}}{|x^{\prime}+\xi^{\prime}|},\frac{x^{\prime}-\xi^{\prime}}{|x^{\prime}-\xi^{\prime}|},T(x,\xi)\right)\end{split}$
where $T(x,\xi)$ is a parameter on the geodesic $e^{tY}(x,\xi)$, we note
$\begin{split}\Phi(\pi_{S}(e^{sU_{j}^{+}}(x,\xi^{1},\cdots,\xi^{d+1})))&=\left(\frac{x^{\prime}+(\xi^{1})^{\prime}}{|x^{\prime}+(\xi^{1})^{\prime}|},\frac{x^{\prime}-(\xi^{1})^{\prime}-2s(\xi^{j+1})^{\prime}+s^{2}(x^{\prime}+(\xi^{1})^{\prime})}{|x^{\prime}-(\xi^{1})^{\prime}-2s(\xi^{j+1})^{\prime}+s^{2}(x^{\prime}+(\xi^{1})^{\prime})|},T\right);\\ \Phi(\pi_{S}(e^{sU_{j}^{-}}(x,\xi^{1},\cdots,\xi^{d+1})))&=\left(\frac{x^{\prime}+(\xi^{1})^{\prime}-2s(\xi^{j+1})^{\prime}+s^{2}(x^{\prime}-(\xi^{1})^{\prime})}{|x^{\prime}+(\xi^{1})^{\prime}-2s(\xi^{j+1})^{\prime}+s^{2}(x^{\prime}-(\xi^{1})^{\prime})|},\frac{x^{\prime}-(\xi^{1})^{\prime}}{|x^{\prime}-(\xi^{1})^{\prime}|},T\right).\end{split}$
We claim the nontrivial component
$\frac{x^{\prime}\mp(\xi^{1})^{\prime}-2s(\xi^{j+1})^{\prime}+s^{2}(x^{\prime}\pm(\xi^{1})^{\prime})}{|x^{\prime}\mp(\xi^{1})^{\prime}-2s(\xi^{j+1})^{\prime}+s^{2}(x^{\prime}\pm(\xi^{1})^{\prime})|}$ (A.4)
gives a circle segment on the sphere ${\mathbb{S}}^{d}$. This is because $x\mp\xi^{1}-2s\xi^{j+1}+s^{2}(x\pm\xi^{1})$ lies in the null cone of the $3$-dimensional space $V_{x,\xi^{1},\xi^{j+1}}$ spanned by $x,\xi^{1}$ and $\xi^{j+1}$, and the claim follows from the fact that the intersection of ${\mathbb{S}}^{d}$ with the two dimensional plane $V_{x,\xi^{1},\xi^{j+1}}\cap\{x_{0}=1\}$ is a circle. It is also clear that the curve (A.4) cannot be a single point (otherwise, $x^{\prime},(\xi^{1})^{\prime},(\xi^{j+1})^{\prime}$ are linearly dependent, which is impossible).
A.2. Spectral gap
We recall the fractal uncertainty principle in the regime of unitarity bound.
Definition A.2.
Let $A$ be a compact set in ${\mathbb{R}}^{d}$. Let $\nu>0$ and $0\leq\alpha_{0}<\alpha_{1}$.
  • We say $A$ is $\nu$-porous from scales $\alpha_{0}$ to $\alpha_{1}$, if for any ball $B(x,r)$ of radius $r\in(\alpha_{0},\alpha_{1})$, there exists a ball $B(x^{\prime},\nu r)$ of radius $\nu r$ inside $B(x,r)$ such that $A\cap B(x^{\prime},\nu r)=\varnothing$.
  • We say $A$ is $\nu$-porous along lines from scales $\alpha_{0}$ to $\alpha_{1}$, if for any line segment $\ell$ of length $r\in(\alpha_{0},\alpha_{1})$, there exists a point $x\in\ell$, such that $A\cap B(x,\nu r)=\varnothing$.
If a set is $\nu$-porous ($\nu$-porous along lines) from scales $\alpha_{0}$ to $\alpha_{1}$ for some $\nu>0$, we will say it is porous (line porous) from scales $\alpha_{0}$ to $\alpha_{1}$.
Bourgain–Dyatlov [BoDy18] establishes the following fractal uncertainty principle (see [JiZh20] for a quantitative version).
Theorem A.3.
Let $\nu>0$ and $h\in(0,1)$. Suppose $A_{h}$ and $B_{h}$ are two compact subsets of $[0,1]$ that are $\nu$-porous from scales $h$ to $1$, then there exist $\beta=\beta(\nu)>0$ and $C=C(\nu)>0$ independent of $h$, such that
$\|\mathbbm{1}_{A_{h}}\mathcal{F}_{h}\mathbbm{1}_{B_{h}}\|_{L^{2}({\mathbb{R}})\to L^{2}({\mathbb{R}})}\leq Ch^{\beta}.$
In a recent breakthrough, Alex Cohen [Co25b] proves a fractal uncertainty principle in higher dimensions (see [JZZ26] for a quantitative version).
Theorem A.4.
Let $v>0$ and $h\in(0,1)$. Suppose $A_{h}$ and $B_{h}$ are two compact subsets of $0,1]^{d}$ that
  • $A_{h}$ is $\nu$-porous from scales $h$ to $1$;
  • $B_{h}$ is $\nu$-porous along lines from scales $h$ to 1.
Then there exist $\beta=\beta(\nu,d)>0$ and $C=C(\nu,d)>0$ independent of $h$, such that
$\|\mathbbm{1}_{A_{h}}\mathcal{F}_{h}\mathbbm{1}_{B_{h}}\|_{L^{2}({\mathbb{R}}^{d})\to L^{2}({\mathbb{R}}^{d})}\leq Ch^{\beta}.$
By [KM25, Proposition 4.14], Cohen’s theorem applies to a general non-degenerate phase. By [DyZa16] we have the following corollary.
Corollary A.5.
[…]
This leaves the question of which limit sets are line porous. We first give some basic properties of line porosity.
Lemma A.6.
Let $A\subset{\mathbb{R}}^{d}$ be a compact set.
  1. (1)
    Suppose $A$ is $\nu$-porous along lines from scales $\alpha_{0}$ to $\alpha_{1}$, and $F:{\mathbb{R}}^{d}\to{\mathbb{R}}^{d}$ is a diffeomorphism such that
    $C_{1}^{-1}|x-y|\leq|F(x)-F(y)|\leq C_{1}|x-y|,\quad|\partial_{i}\partial_{j}F(x)|\leq C_{2}.$
    Suppose $0<\alpha_{0}<\alpha_{1}<\frac{\nu}{C_{1}C_{2}d}$. Then $F(A)$ is $\frac{\nu}{2C_{1}^{2}}$-porous along lines from scales $C_{1}\alpha_{0}$ to $C_{1}\alpha_{1}$.
  2. (2)
    Suppose there exists $\nu>0$, such that any line segment $\ell$ of length $r\in(\alpha_{0},\alpha_{1})$ with midpoint in $A$ contains a point $x\in\ell$ such that $B(x,\nu r)\cap A=\varnothing$. Then $A$ is $\nu/2$-porous along lines from scales $\alpha_{0}$ to $\alpha_{1}$.
Proof.
  1. (1)
    This is [KM25, Lemma 4.10].
  2. (2)
    If $A$ is not $\nu/2$-porous from scales $\alpha_{0}$ to $\alpha_{1}$, then there exists a line segment $\ell$ with length $r\in(\alpha_{0},\alpha_{1})$, such that the middle point $x$ of $\ell$ satisfies $B(x,\nu r/2)\cap A\neq\varnothing$. Suppose $x^{\prime}\in B(x,\nu r/2)\cap A$. Let $\ell^{\prime}$ be the line segment with midpoint $x^{\prime}$ parallel to $\ell$ and of length $r$. Then for any $y\in\ell^{\prime}$, we have $B(y,\nu r)\cap A\neq\varnothing$, which contradicts the assumption. ∎
In particular, by Lemma A.6 (1), it makes sense to talk about line porosity on a manifold. We provide the following characterization for line porosity of the limit set $\Lambda_{\Gamma}\subset{\mathbb{S}}^{d}$.
Proposition A.7.
The limit set $\Lambda_{\Gamma}\subset{\mathbb{S}}^{d}$ is line porous from scales 0 to $1$ if and only if it does not contain a round circle.
Proof.
If the limit set $\Lambda_{\Gamma}$ contains a circle, then it is not line porous because a circle is not line porous. We would like to show the other direction, i.e., if $\Lambda_{\Gamma}$ is not line porous, then it must contain a circle.
Let $K_{S}\subset F^{*}X$ be the trapped set in the frame bundle. First, we show that the following assumption on the frame bundle implies the line porosity of $\Lambda_{\Gamma}$: there exists a constant $T>0$ such that for any $z\in K_{S}$, there exists $s\in[-T,T]$ such that
$e^{sU_{1}^{+}}z\notin K_{S}.$ (A.5)
By compactness of $K_{S}$, this implies that there exists a uniform constant $R>0$ such that for some $s_{1}\in[-T,T]$, we have $B(e^{s_{1}U_{1}^{+}}z,R)\cap K_{S}=\varnothing$. For any point $z_{0}\in K_{S}$ and a flow line $\ell_{z_{0},s_{0}}=\{e^{sU_{1}^{+}}z_{0}:-s_{0}\leq s\leq s_{0}\}$ with $s_{0}\in(0,1)$, we may choose $t_{0}\in{\mathbb{R}}$ such that $e^{t_{0}}s_{0}=T$. Since $e^{tY}\circ e^{sU_{1}^{+}}(z)=e^{e^{-t}sU_{1}^{+}}\circ e^{tY}(z)$ (note the composition is different from matrix multiplication since the action is on the right), we have
$e^{-t_{0}Y}(\ell_{z_{0},s_{0}})=\ell_{e^{-t_{0}Y}(z_{0}),e^{t_{0}}s_{0}}=\ell_{e^{-t_{0}Y}(z_{0}),T}.$
By our assumption (A.5), there exists $s_{1}\in[-T,T]$ such that
$B\left(e^{s_{1}U_{1}^{+}}\left(e^{-t_{0}Y}(z_{0})\right),R\right)\cap K_{S}=\varnothing.$
In particular, for any $U^{+}=\sum a_{j}U_{j}^{+}$ and $|a_{j}|\leq R^{\prime}$ for some $R^{\prime}=R^{\prime}(R)>0$, we have
$e^{-t_{0}Y}\left(e^{e^{-t_{0}}U^{+}}\circ e^{e^{-t_{0}}s_{1}U_{1}^{+}}(z_{0})\right)=e^{U^{+}}\circ e^{s_{1}U_{1}^{+}}\left(e^{-t_{0}Y}(z_{0})\right)\notin K_{S}.$
In other words, $z_{1}=e^{e^{-t_{0}}s_{1}U_{1}^{+}}(z_{0})\in\ell_{z_{0},s_{0}}$ (note $e^{-t_{0}}s_{1}\in[-s_{0},s_{0}]$) satisfies
$e^{\sum a_{j}U_{j}^{+}}(z_{1})\notin K_{S},\quad|a_{j}|\leq s_{0}R^{\prime}T^{-1}.$ (A.6)
Now to show $\Lambda_{\Gamma}$ is line porous, we take $x\in\Lambda_{\Gamma}$. Pick another point $x^{\prime}\in\Lambda_{\Gamma}$ and consider $y\in S^{*}\mathbb{H}^{d+1}$ that corresponds to $(x^{\prime},x)$ under the Hopf parametrization. More precisely, the geodesic emanating from $y$ converges to $x^{\prime}$ in the forward direction and converges to $x$ in the backward direction. Let $y_{0}\in S^{*}X$ be the projection of $y$ to the quotient. Let $z_{0}\in F^{*}X$ be a lift of $y_{0}$ in the frame bundle and $\ell_{z_{0},s_{0}}$ be a path going through $z_{0}$. Since $z_{0}\in K_{S}$, the above reasoning shows that there is a point $z_{1}\in\ell_{z_{0},s_{0}}$ satisfying (A.6). Projecting this back to a neighbourhood of $x$ shows that a circle segment corresponding to $\ell_{z_{0},s_{0}}$ contains a point such that its neighbourhood of size $\simeq s_{0}$ does not intersect $\Lambda_{\Gamma}$. Note by changing the choice of $x^{\prime}$ and the lift $z_{0}$ in the frame bundle, this covers different circles (in particular, circles in different directions, see (A.4)) passing though the point $x$. Since the notion of line porosity is invariant under local uniform change of coordinates (see Lemma A.6), we conclude that $\Lambda_{\Gamma}$ is line porous from scales 0 to $1$.
Finally, we show that if (A.5) does not hold, then $\Lambda_{\Gamma}$ must contain a circle. Suppose (A.5) does not hold, then there exist $z_{n}\in K_{S}$ and $T_{n}\to\infty$ such that
$e^{sU_{1}^{+}}z_{n}\in K_{S},\quad s\in[-T_{n},T_{n}].$
Since $K_{S}$ is compact, there exists a limit point $z\in K_{S}$ of $\{z_{n}\}$, such that
$e^{sU_{1}^{+}}z\in K_{S},\quad s\in{\mathbb{R}}.$
Projecting this to ${\mathbb{S}}^{d}$ gives a circle (see the discussion at the end of §A.1). ∎
Combining Corollary A.5 and Proposition A.7, we prove Theorem A.1.